write the equation in standard form for the parabola with vertex (0,0) and directrix y=14
Answer by lwsshak3(11628)

write the equation in standard form for the parabola with vertex (0,0) and directrix y=14
**
This is a parabola that opens downwards.
Its standard form of equation: (x-h)^2=-4p(y-k)^2, with (h,k)=(x,y) coordinates of the vertex
For given parabola:
vertex: (0,0) (given)
axis of symmetry=y-axis or x=0
p=14 (distance from directrix to vertex on the axis of symmetry)
4p=56
Equation:
x^2=-56y